# LISTN: Lexicon induction with socio-temporal nuance

### v1

@@ 4.3 Existing lexicons
IncelWiki [^1] https://incels.wiki/w/Incel_Glossary 189
@@
Hatebase [^2] https://hatebase.org/search_results/gender_id%3D6%7Clanguage_id%3Deng 29
@@
Table 1: Pre-existing lexicons used in the evaluation, and the number of terms from each lexicon that occur in the training data.
For the Farrell et al. (2019) dataset, we follow Ribeiro et al. (2021) in excluding the lexicon categories that are not specific to the manosphere (e.g. racism). The Hatebase and IncelWiki lexica are crawled by us and will be released to the community. Combined, this provides a set of 483 terms.
4.4 Evaluation
To evaluate our methods, we first perform a preliminary scoring step that is used to construct a test set. This test set is used as gold standard in our main evaluation.
@@ 4.4 Evaluation
In this step, we evaluate each of the approaches in Sections 3 & 4.2 against the lexicons in Section 4.3. Terms in the lexicons are assigned a True label and all others are considered False. An issue with this approach, however, is that the lexicons suffer from low recall (a known issue; see Lucy and Bamman, 2021), meaning that true in-group terms identified by the methods are often evaluated as false positives. For this reason, we use this scoring only to identify terms to include in our test set (described below). The average precision and area under the receiver-operator characteristic curve of each method are provided in Appendix A.
Test set construction
We retrieve the top-scoring 1000 terms produced by the best baseline and best LISTN approaches from the preliminary scoring step. These terms are manually evaluated by an social psychology PhD student who is a specialist on the manosphere, as well as an author of this paper. Annotators are tasked with assigning a binary label to each word, based on whether or not it strongly signals membership of the manosphere. They are also provided with three example usages of each word and the name of the subreddit where it was observed. Following the expert annotator’s advice, we also annotate “cultural exports” – terms that originated within the manosphere but have spread to be used more widely. In our evaluation, we treat these terms as positives. We further include the terms from the existing lexicons in Table 1 as positive samples.
The resulting test set consists of 1803 words with binary labels, with an inter-annotator agreement score of 0.726 using Cohen’s Kappa, indicating substantial agreement. In cases of disagreement, we revert to the label from the expert annotator. The test set is reasonably balanced, consisting of 944 positive and 859 negative samples. This dataset will be released to the public.
Main evaluation
This dataset is used for evaluation in all experiments in this paper. We report the average precision (AP) and area under the receiver-operator characteristic curve (AUROC) for each method. The AP provides a measure of the model’s precision over all possible recall values; that is, how well the model avoids false positives. AUROC is a measure of its ability to distinguish between positive and negative cases. Davis and Goadrich (2006) show that a model dominates in AUROC space if and only if it dominates in AP space, whereas the converse is not true; as such, we prioritise the AP. We use the randomised permutation test with Monte Carlo approximation ($R=9999$ and $\alpha=0.05$) for significance testing.
@@ 5 Results
Results are shown in Table 2. The best LISTN-CA and LISTN-C methods both outperform the best baseline method (statistically significant, $P<<0.05$). Using LISTN-C produces a higher AP score compared to LISTN-CA for 6 out of the 8 induction methods, including the top-performing method (not statistically significant; $P=0.723$). This is a surprising result, as it indicates that the induction does not benefit from the added author adjacency information (i.e. shared thread interaction) in LISTN-CA. A benefit of this outcome is that the LISTN-C model has fewer parameters and is more computationally efficient to train. The LISTN-C model does still contain social information, in that the source matrix is a user-content representation, meaning that users are represented as being similar if they use the same words. By comparison, the standard word2vec matrix factorisation formulation only considers linguistic context – words are similar if they co-occur with the same context words (Goldberg and Levy, 2014).
@@
For both LISTN-CA and LISTN-C, the 5-cluster approach performs the best. Notably, it performs better than category-level induction, which computes centroids of a similar level of granularity, indicating that the embedding clustering is identifying useful sub-groups which are not evident from the subreddit structure. Using more clusters ($K=20$ and $K=100$) does not improve performance. A possible explanation might be that smaller clusters lead to a less precise or less informative centroid which overemphasises irrelevant words. Subreddit-level induction (a more granular version of the category-level method) performs the worst of the centroid-based approaches for both models and metrics. However, averaging over the full community performs the second worst, indicating that it is necessary to take into account some level of sub-group linguistic specialisation.
@@
All baseline approaches outperform the random baselines. The worst-performing baseline is the word2vec-based method. For both LISTN models, the embedding expansion approaches (“bootstrap”) perform the worst as well. A possible reason for this might be that the seed lexicons are not varied enough, such that they are missing some area of the embedding space. The NPMI baseline performs the best on AUROC, whereas the category-level NPMI performs the best for the AP. Generally, the two measures tend to co-vary. In all cases, the NPMI outperforms the PPMI approaches, which corresponds with the results of Lucy and Bamman (2021). Interestingly, using month-level NPMI scoring does not improve over the time-aggregated NPMI, indicating that naive models for accounting for time are not sufficient.
6 Analysis
Beyond the improved performance of lexicon induction methods (as illustrated in Section 5), an advantage of this approach is that it produces temporal embeddings for words and for users. In this section, we use these embeddings to investigate changes in words over time and sub-group specialisation of in-group language in the manosphere.
@@ 6.1 Change in word representations
Figure 1: The relationship between word frequency and embedding stability .
@@
Recall that the training objective includes a regularisation term that penalises changes between successive word representations, meaning that changes in embeddings are only permitted if they reduce the reconstruction error sufficiently. In other words, word embeddings change as much as is needed to represent the data, but no more. To determine how much a word changes between two successive timesteps, we use the cosine distance between the embeddings. Its sum over all timesteps is calculated as a measure for how much the word representation varies over the training window.
@@
Using this metric, we find that the words with the lowest variation are stopwords, e.g. from, here, there. This makes sense as the embeddings are a function of how a word is used throughout a community, and these words are not expected to be impacted by the social evolution of the group. At the opposite end of the spectrum, we observe less common words, including betta, lackeys, wooo, fragment, thirstiest, and booklet. Noting that there seems to be a relation between word frequency and embedding variation, we calculate their Spearman correlation and find a strongly negative correlation, $\rho=-0.77$ ($P<<0.05$).
@@
Figure 1 shows the relation between the log frequency of a word and its cumulative embedding variation, for all words in the training data (grey dots). The strong negative correlation only appears to hold for frequencies under approx. 10 000; beyond this point, changes become close to zero. We also observe that the variance in the amount of change covaries with the log frequency, such that a large range of values is observed for low-frequency words. We have two hypotheses for why this might be. Firstly, it may simply be a computational issue, whereby words that are less common do not have high-quality representations due to insufficient training data, and therefore their embeddings are less stable. The second option is that words that are less common truly have more pliability in their colloquial usage, and are used in more subtle, differentiated ways. We expect the reality is a combination of the two.
@@
To investigate this idea, we also plot the confirmed in-group language along the same axes (colourful dots) in Figure 1. Dotted lines indicate straight-line regressions to the whole population and to the in-group language sub-population; we observe that the regression to the in-group population is less steep. This seems to suggest that in-group language terms are more stable and less pliable in how they are used within this community – providing some support for the idea that the variation is only partly explained by frequency. The different colours represent the predictions made by the best models in Section 5, showing the top 1000 words predicted by each approach. The two methods appear to favour different areas in the frequency distribution, with the NPMI-based model being biased to the lower end of the spectrum, and the LISTN-based model to the mid and higher side. Future work may investigate combining these methods for lexicon induction, as they appear to have different strengths.
@@
Further investigating the low-frequency, low-variation terms, we find that it includes several mentions of medications (e.g. lamictal, pristiq, seroquel). These are concepts with very specific, technical meanings that are not likely to vary due to socio-temporal changes in the community. We also note the term thotaudit, which refers to the hashtag of a harassment campaign targeting sex workers [^3] https://www.vice.com/en/article/thotaudit-databases-of-sex-workers-and-reporting-them-to-paypal/. On the other end of the spectrum, the terms submission and bimbo display a large amount of variation relative to their frequency. A possible reason for this might be that these terms occupy different niches for the subcommunities of the manosphere, or that they underwent a socio-semantic change during the training period. While more exploration of this embedding space is warranted, it is evident that it captures some interesting and useful qualities of the data.
6.2 Correlation with violent language
Figure 2: Relevance of violence lexicon and in-group lexicon to different subreddits.
In this section, we explore how violence interacts with in-group language, using the lexicon of confirmed manosphere language from Section 5. Chancellor et al. (2016) found that lexical variation is associated with more extreme content in pro-anorexia communities, but this has not been explored in the context of the manosphere.
Similar to De Kock (2024), we calculate centroids for each of the 50 subreddits in the dataset as the mean of all users who participated in that subreddit at a given timestep. Then, the relevance score of the lexicon is given by the dot product of the lexicon centroid with each of the sub-community centroids. For each subreddit, a score is calculated at each timestep, since the community and word embeddings are both dynamic. Figure 2 shows the relevance scores for the manosphere lexicon (top) and the violence lexicon (bottom). Community sizes are reflected by the relative sizes of the different dots.
Based on this characterisation, Incel subreddits display the highest relevance scores for the manosphere lexicon, with the exception of the r/foreveralonedating and r/foreveralone subreddits, who also have the lowest violence scores of the incel groups. Although the taxonomy of Ribeiro et al. (2021) placed this subreddit in the Incels category, Bachaud (2024) states r/foreveralone actively distances itself from Incels, which is supported by our analysis.
The Pick-up Artists, who have relatively low violence scores, also have relatively low in-group language scores. On the other hand, the MGTOW and MRA groups have higher violence scores but relatively low in-group language scores, showing that these metrics are not always correlated. Calculating the Spearman correlation between the violence and in-group language scores of all users in the dataset, we find a moderate positive correlation of $\rho=0.53$, $(P<<0.05)$. Breaking this down per manosphere category, we observe the highest correlation ($\rho=0.65$) for the Incels, and the lowest for MRA ($\rho=0.45$).
The relevance of the manosphere lexicon is the lowest for the mental health subreddits, which is consistent with our expectations for the control group. The r/suicidewatch group has an above average violence score, whereas the r/depression subreddit is on par with the community average.
@@ 6.3 Sub-group linguistic specialisation
The analysis in Section 6.2 indicates that there is a complex relationship between the utilisation of in-group language (as defined by our lexicon) and violent language, which differs at the sub-group level. This, again, points to the existence of sub-group linguistic specialisation among manosphere groups, as mentioned in Section 1. In this section, we further investigate this idea. The hypothesis is as follows: if manosphere sub-groups share a similar vernacular but express it to varying degrees, we would expect there to be a (nonlinear) correlation between their utilisation of different words in the lexicon. On the other hand, if the groups favour disjoint parts of the lexicon, there would be no such correlation.
To evaluate this idea, we investigate the relevance of individual terms in the lexicon to each of the subcommunities. Table 3 shows the highest-ranking lexicon terms for each sub-community. These words align with the known focuses for the different groups; for example, the Pick-up Artists use terms such as IOI (indicators of interest) and SMV (sexual market value) and LTR (long-term relationship). To enable future reseach on these communities, we will publish the score of each term in the lexicon for each sub-community along with the lexicon.
The Spearman correlations of the word relevance across different subcommunities are shown in Figure 3. The Incels have negative correlations with all other communities (with a minimum $\rho=-0.240$ compared to the Men’s Rights Activists), showing that they are distinctive in their in-group language.
The largest correlation ($\rho=0.729$) is observed between the Pick-up Artists (PuA) and The Red Pill (TRP). This has been observed in prior work in sociology: Bachaud (2024) states that PuA terminology is plentiful in the TRP community, as they have a shared focus on seduction. The MGTOWs have a fairly strong correlation with TRP ($\rho=0.615$) and MRA ($\rho=0.654$). This aligns with their historical connections: Bachaud (2024) states that MGTOW was created by a group of MRAs and that their ideology is “close to both Red Pillers (critical about female nature) and MRAs (denouncing feminism and structural biases against men)”. A much weaker correlation is observed between MGTOW and PuA ($\rho=0.282$), which can be explained by their conflicting ideologies: PuAs are intent on pursuing women, whereas the main premise of MGTOW is to eschew contact with women. This captures an interesting insight into the relationship between PuAs, MGTOWs and TRPs: non-overlapping parts of the TRP interests are shared by both other groups.
The Criticism groups, who are reformed manosphere members, also correlate positively with MGTOW, MRA, TRP and PuA. Investigating their most relevant words (Table 3), we note that it mostly lists references to the manosphere itself.
@@
Table 3: Highest relevance lexicon terms per category. Words in bold are novel contributions, i.e. not included in the lexicons in Table 1.
7 Related work
A number of interesting prior works relate to ours. We highlight some of their findings here.
Chancellor et al. (2016) investigate linguistic variation and the processes by which existing words are modified to create neologisms in the context of pro-anorexia groups, finding that lexical innovations are associated with increased engagement and more extreme content in the context of pro-anorexia online communities. Danescu-Niculescu-Mizil et al. (2013) explored the role of linguistic evolution in groups, using snapshot n-gram language models to measure the deviation of a user from the community. They illustrate that the non-adoption of lexical innovations is predictive of user churn in online communities. Zhang et al. (2017) characterises how distinctive and temporally dynamic online communities are in their language, finding that communities with more dynamic and distinctive language are likely to have higher retention of members.
Finally, Stewart and Eisenstein (2018) and Del Tredici and Fernández (2018) investigate factors that influence the dissemination of new lexical innovations, finding that linguistic and social factors both play a role in this process.
@@ 8 Conclusion
In this paper, we introduced and evaluated a novel framework for lexicon induction based on pretrained word embeddings. We compare several approaches based on two different basic models to methods from prior work, finding that the proposed methods achieve better scores on a newly crafted, expert-validated test set. A set of 455 new manosphere terms are to be released based on this analysis, with relevance scores for each manosphere sub-group.
Our work further provides novel insights into in-group language, including its embedding stability and correlation to violent language. The investigation into sub-group linguistic specialisation provides insight into the relationships between the manosphere subcommunities, which is supported by claims from prior work in sociology.
@@
The experiments in this paper make use of data developed in prior work. Ribeiro et al. (2021) state that the ethical standard guidelines of Rivers and Lewis (2014) were followed in the construction of this dataset, which includes not making any attempt to de-anonymize or link users across platforms. Even so, we recognise that the users did not consent to having their data analysed for the purpose of scientific experiments. However, given the real and current threat posed by this group, we view it is an acceptable use of the data. To limit the exposure of individuals, we do not provide any quotations or usernames and all analyses are performed at an aggregate level.
As we are mindful of potential mental health repercussions of this work to collaborators, participants (e.g. annotators) who were exposed to texts originating from manosphere communities were encouraged to attend a support group for researchers who work with extreme content.
@@
There are a number of simplifying assumptions that are made in this work. All models evaluated (LISTN and baselines) provide only one score per word, and hence do not account for sense variation or polysemy. We further include only single-word lexical variations in our evaluation and exclude non-alphanumeric characters. Our motivation for these choices is provided in Section 4.1.
The LISTN and word2vec models are trained on data from 2018. Given the dynamic nature of in-group language, this limits the applicability of the lexicon. Given recent limitations placed on Reddit data access, the Ribeiro et al. (2021) dataset still constitutes one of the largest and most varied existing manosphere datasets. Future work may look at applying these methods to more modern datasets (possibly in the context of specialised forums like incels.is); however, the evaluations in this work serve to validate our lexicon induction approach.
Appendix A Preliminary evaluation

### v3

@@ 6 Test set construction
IncelWiki 189
@@
Hatebase 29
@@
Table 1: Existing lexicons used in the evaluation.
@@
In this step, we evaluate each of the approaches in Sections 4 and 5.1 against the lexicons in Table 1. Terms in the lexicons are assigned a positive label; all others are considered negatives. The average precision and area under the receiver-operator characteristic curve of each method are provided in Appendix B. An issue with this approach, however, is that the lexicons suffer from low recall (a known issue; see Lucy and Bamman, 2021), meaning that true in-group terms identified by the methods are often wrongly evaluated as false positives. For this reason, we use this scoring only to identify words to include in our test set.
Manual validation
We retrieve the top-scoring 1000 terms produced by the best baseline and best LISTN approaches from the preliminary scoring step. These terms are manually evaluated by the author of this paper as well as a social psychology PhD student who is a specialist on the manosphere. Annotators are tasked with assigning a binary label to each word, based on whether or not it strongly signals membership of the manosphere. Annotators are also provided with three example usages of each word and the corresponding subreddit name. Following the expert annotator’s advice, we also annotate “cultural exports” – terms that originated within the manosphere but have spread to be used more widely. In our evaluation, we treat these terms as positives. We further include the terms from the existing lexicons in Table 1 as positive samples.
The resulting test set consists of 1803 words with binary labels, with an inter-annotator agreement score of 0.726 using Cohen’s Kappa, indicating substantial agreement. In cases of disagreement, we use the label from the expert annotator. The test set is reasonably balanced, consisting of 944 positive and 859 negative samples.
@@ 7 Results
Results are shown in Table 2. The best LISTN-CA and LISTN-C methods both outperform the best baseline method (statistically significant, $P<<0.05$). Using LISTN-C produces a higher AP score compared to LISTN-CA for 6 out of the 8 induction methods, including the top-performing method (not statistically significant; $P=0.723$). This is a surprising result, as it indicates that the induction does not benefit from the added author adjacency information (i.e. shared thread interaction) in LISTN-CA. A benefit of this outcome is that the LISTN-C model has fewer parameters and is more computationally efficient to train. The LISTN-C model does still contain social information, in that the source matrix is a user-content representation, meaning that users are represented as being similar if they use the same words.
@@
For both LISTN-CA and LISTN-C, the 5-cluster approach performs the best ($AP=0.75$ and $0.77$ respectively). Using a larger number of clusters ($K=20$ and $K=100$) leads to worse performance. A possible explanation might be that smaller clusters lead to less meaningful centroids that overemphasise irrelevant words. On the other hand, averaging over the full community performs poorly, indicating that it is necessary to take into account some level of sub-group linguistic specialisation. This is corroborated by the relatively low scores of the bias and bootstrap methods, which also do not account for community structure.
Notably, the 5-cluster method performs better than category-level induction, which computes centroids of a similar level of granularity. Subreddit-level induction (a more granular version of the category-level method, with 52 subgroups) performs the worst of the centroid-based approaches for both models and metrics. This is an unexpected result; one would expect this approach to perform on-par with the other centroid-based approaches. The cluster-based methods are intended to capture social structure at the intra-subreddit and inter-subreddit level, and assign each user to one cluster only. The subreddit representations average over all users who have contributed to a given subreddit per timestep, meaning that one user may form a part of many representations. This might include users who are incidental posters, which would add noise to the centroids. This effect would be more pronounced at the subreddit level than the category level, as there is less inter-category activity, which would explain why the category aggregation is better than subreddit aggregation but still worse than the clustering-based methods. Using a user weighting based on post frequency might resolve this issue. However, since the subreddit structure does not translate to other platforms, it is beneficial that the best method does not require it.
@@
All baseline approaches outperform the random baseline. The worst-performing baseline is the word2vec-based method. For both LISTN models, the lexicon expansion approaches (“bootstrap”) perform the worst as well. A possible reason for this might be that the seed lexicons are not varied enough, such that they are missing some area of the embedding space. The NPMI baseline performs the best on AUROC, whereas the category-level NPMI performs the best for the AP. In all cases, NPMI outperforms the PPMI, which corresponds with the results of Lucy and Bamman (2021). Interestingly, using month-level NPMI scoring does not improve over the time-aggregated NPMI, indicating that these superficial methods for accounting for time are not sufficient.
8 Embedding analysis
Beyond the improved performance over prior lexicon induction methods, an advantage of this approach is that it produces dynamic embeddings for words and users, in the same space. In this section, we use these embeddings to investigate (i) changes in words over time and (ii) sub-group linguistic specialisation in the manosphere.
@@ 8.1 Change in word representations
Figure 1: Word frequency versus embedding stability.
@@
Recall that the training objective includes a regularisation term that penalises large changes between temporally successive word representations, meaning that changes in embeddings are only permitted if they reduce the reconstruction error sufficiently. In other words, word embeddings change as much as is needed, but no more. To determine how much a word changes between two successive timesteps, we calculate the cosine distance between successive $W_{t}$ representations. The sum of these distances all timesteps is referred to as the cumulative embedding variation (CEV).
@@
We find that the words with the lowest CEV are stopwords, e.g. from, here, there. This makes sense, given that the embeddings are a function of how a word is used throughout a community, and these words are not expected to be impacted by the social evolution of the group. At the opposite end of the spectrum, showing the most variation, we observe less common words such as lackeys, wooo, fragment, and booklet. Noting that there seems to be a relation between word frequency and CEV, we calculate their Spearman correlation, yielding strongly negative value, $\rho=-0.77$ ($P<<0.05$).
@@
Figure 1 shows the relation between the log frequency of a word and CEV, for all words in the training data (grey dots). We also plot the confirmed in-group language along the same axes (colourful dots) in Figure 1. The strong negative correlation only appears to hold for frequencies under approx. 10 000; beyond this point, changes become close to zero. We have two hypotheses for why this might be. Firstly, it may simply be a computational issue, whereby words that are less common do not have high-quality representations due to insufficient training data, and therefore their embeddings are less stable. The second option is that words that are less common have more pliability in their colloquial usage. We expect the reality is a combination of the two.
We further observe that the variance in CEV covaries with frequency, such that a larger range of values is observed for low-frequency words. The confirmed in-group language tend to exhibit lower CEV per frequency, relative to the general vocabulary. Dotted lines indicate straight-line regressions to CEV of the whole vocabulary (black) and to the in-group language lexicon (yellow). We observe that the gradient for the in-group language is indeed less steep. This suggests that in-group language terms are more stable and less pliable in how they are used within this community, relative to what we would expect given their frequency. This could be a result of their special status as a social artefact in the group: there is an incentive to use the term “correctly” to signal group belonging.
@@
The different colours represent the predictions made by the best models in Section 7, showing the top 1000 words predicted by the NPMI-category and LISTN-C (cluster-5) methods. The two methods appear to favour different areas in the frequency distribution, with the NPMI-based model being biased to the lower end of the spectrum, and the LISTN-based model to the mid and higher side. Future work may investigate combining these methods for lexicon induction.
@@
Investigating the low-frequency, low-variation words, we find that it includes several mentions of medications (e.g. lamictal, pristiq, seroquel). These are concepts with specific technical meanings that are not likely to vary due to socio-temporal changes in the community. We also note the term thotaudit, which is the hashtag of a harassment campaign targeting sex workers. On the other end of the spectrum, submission and bimbo display a large amount of variation relative to their frequency. A possible reason for this might be that these terms occupy different niches for the subcommunities of the manosphere, or that they underwent a socio-semantic change during the training period. While more exploration of this embedding space is warranted, it is evident that it captures some interesting and useful qualities of the data.
@@ 8.2 Sub-group linguistic specialisation
The results in Section 7 show that lexicon induction benefits from accounting for social sub-group structure, suggesting that there might be sub-group linguistic specialisation. In this section, we further investigate this idea. The hypothesis is: if manosphere categories share a similar vernacular but express it to varying degrees, we would expect there to be a (nonlinear) correlation between their preference for different words in the lexicon. If the groups favour disjoint parts of the lexicon, there would be no such correlation.
@@
Table 3: Highest relevance lexicon terms per category.
To evaluate this idea, we firstly calculate the relevance of individual words in the lexicon to each of the manosphere categories. Table 3 shows the highest-ranking lexicon words for each grouping. These words align with their known thematic interests; for example, the Pick-up Artists use terms such as IOI (“indicators of interest”) and SMV (“sexual market value”). To enable future research on these communities, these scores are also published.
We then find the Spearman correlation between the per-word relevance scores for each pair of communities, shown in Figure 2. We observe that they correlate to varying degrees, with the largest correlation ($\rho=0.729$) between the Pick-up Artists (PuA) and The Red Pill (TRP). This has been observed in prior work in sociology: Bachaud (2024) states that PuA terminology is plentiful in the TRP community, as they have a shared focus on seduction. The MGTOWs have a fairly strong correlation with TRP ($\rho=0.615$) and MRA ($\rho=0.654$). This aligns with their historical connections: Bachaud (2024) states that MGTOW was created by a group of MRAs and that their ideology is “close to both Red Pillers (critical about female nature) and MRAs (denouncing feminism and structural biases against men)”. A much weaker correlation is observed between MGTOW and PuA ($\rho=0.282$), which can be explained by their conflicting ideologies: PuAs are intent on pursuing women, whereas the main premise of MGTOW is to eschew contact with women. This captures an interesting insight into the relationship between PuAs, MGTOWs and TRPs: non-overlapping parts of the TRP interests are shared by both other groups.
The Incels have negative correlations with all other communities (with a minimum $\rho=-0.240$ compared to the Men’s Rights Activists), showing that they have the most specialised in-group language. This, too, is supported by manosphere history: while the other manosphere groups have fairly intertwined origin stories, the Incels emerged from a separate line of new online platforms such as LoveShy (whose administrators openly supported mass killers and femicide) with cross-pollination from 4/chan and Reddit (Bachaud, 2024).
The Criticism groups, who are reformed manosphere members, correlate positively with MGTOW, MRA, TRP and PuA. Investigating their most relevant words (Table 3), we note that it mostly lists references to the manosphere itself.
@@ 9 Conclusion
In this paper, we introduced a novel framework for lexicon induction based on dynamic word and user embeddings. We compare several approaches based on two different basic models to methods from prior work, finding that the proposed methods achieve better scores on a newly crafted, expert-validated test set. A set of 455 new manosphere terms are to be released based on this analysis, with relevance scores for each manosphere sub-group. Our work provides novel insights into in-group language, including its embedding stability. The investigation into sub-group linguistic specialisation provides insights into the relationships between the manosphere subcommunities, which corroborate claims from prior work in sociology.
@@
The experiments in this paper make use of data developed in prior work. Ribeiro et al. (2021) state that the ethical standard guidelines of Rivers and Lewis (2014) were followed in the construction of this dataset, which includes not making any attempt to de-anonymize or link users across platforms. Even so, we recognise that the users did not consent to having their data analysed for the purpose of scientific experiments. However, given the real and current threat posed by this group, we view it as an acceptable use of the data. To limit the exposure of individuals, we do not provide any quotations or usernames, and all analyses are performed at an aggregate level.
As we are mindful of potential mental health repercussions of this work to collaborators, participants (including annotators) who were exposed to texts originating from manosphere communities were encouraged to attend a support group for researchers who work with extreme content.
@@
There are a number of simplifying assumptions that are made in this work. All models evaluated (LISTN and baselines) provide only one score per word, and hence do not account for sense variation or polysemy. We further include only single-word lexical variations in our evaluation and exclude non-alphanumeric characters. Our motivation for these choices is provided in Section 5.
The LISTN and word2vec models are trained on data from 2018. Given the dynamic nature of in-group language, this limits the applicability of the lexicon to current data; however, the evaluations in this work serve to validate our lexicon induction approach. Given recent limitations placed on Reddit data access, the Ribeiro et al. (2021) dataset still constitutes one of the largest and most varied existing manosphere datasets. Future work may look at applying these methods to more modern datasets (possibly in the context of specialised forums like incels.is).
Acknowledgements
This work is supported by the Hallmark Research Initiative on Fighting Harmful Online Communication. The author would like to thank Macken Murphy, Ed Hovy, and Indigo Orton for their valuable inputs.
Appendix A Hyperparameters
Hyperparameters for the representation learning component are shown in Table 4.
Parameter Value
Early stopping patience 5 epochs
Early stopping tolerance 0.001
K 100
Batch size 80
Weight of A 1
Weight of C 1
Learning rate 0.001
$\lambda_{1}$ 1
$\lambda_{2}$ 1
$c_{0}$ scaler 0.01
Max epochs 100
Table 4: Hyperparameters for training the Cerberus system.
Appendix B Preliminary evaluation